OpenStudy (anonymous):

A rectangle has a width w that is 5 units longer than its length l. Which equation expresses the rectangle's area, A?

OpenStudy (jango_in_dtown):

here w=5+l so area A=lw=l(5+l) that is A=l^2+5l

OpenStudy (anonymous):

Thanks. what about If the area of a rectangle with width x can be represented with the expression A(x) = x(14 – x), what is the perimeter of the rectangle?

OpenStudy (jango_in_dtown):

area=length * width, given A=x(14-x) since x is width, so length is 14-x perimeter=2(length+width)=2(14-x+x)=2*14=28